# EnScale: Temporally-consistent multivariate generative downscaling via proper scoring rules

### v2

@@ Abstract
The practical use of future climate projections from global circulation models (GCMs) is often limited by their coarse spatial resolution, requiring downscaling to generate high-resolution data. Regional climate models (RCMs) provide this refinement, but are computationally expensive. To address this issue, machine learning models can learn the downscaling function, mapping coarse GCM outputs to high-resolution fields. Among these, generative approaches aim to capture the full conditional distribution of RCM data given coarse-scale GCM data, which is characterized by large variability and thus challenging to model accurately. We introduce EnScale, a generative machine learning framework emulating the full GCM-to-RCM map by training on multiple pairs of GCM and corresponding RCM data. It first adjusts large-scale mismatches between GCM and coarsened RCM data, followed by a super-resolution step to generate high-resolution fields. To efficiently model the high-dimensional output, the super-resolution step employs a novel class of sparse local stochastic layers. Both steps employ generative models optimized with the energy score, a proper scoring rule. Compared to state-of-the-art ML downscaling approaches, our setup reduces computational cost by about one order of magnitude. EnScale jointly emulates multiple variables – temperature, precipitation, solar radiation, and wind – spatially consistent over Central Europe. In addition, we propose a variant EnScale-t that enables temporally consistent downscaling. We establish a comprehensive evaluation framework across various categories including calibration, spatial structure, extremes, and multivariate dependencies. Comparison with diverse benchmarks demonstrates EnScale’s strong performance and computational efficiency. EnScale offers a promising approach for accurate and temporally consistent RCM emulation.
@@ 1 Introduction
General circulation models (GCMs) provide us with future climate projections that enable assessing potential changes in climate under different greenhouse gas emission scenarios. Multiple research groups around the world develop and maintain GCMs that incorporate different modeling assumptions and parameterizations to simulate the climate system of the Earth. By virtue of international efforts such as the Coupled Model Intercomparison Project (CMIP) [Eyring2016OverviewOrganization, Taylor2012AnDesign], large ensembles of GCM data are available for various Representative Concentration Pathway (RCP) / Shared Socioeconomic Pathway (SSP) scenarios.
@@
However, these physics-based models are typically computed at a coarse spatial resolution. GCMs are able to simulate large-scale climate patterns, but often lack the spatial resolution needed to capture climate variations at impact-relevant regional and local scales [Benestad2010DownscalingExtremes, Maraun2016BiasReview, Rampal2024EnhancingLearning, Wilby1997DownscalingLimitations]. Downscaling describes the process of increasing the spatial and/or temporal resolution of GCM outputs. The common approach of dynamical downscaling relies on process-based regional climate models (RCMs) that use GCM data as boundary conditions, i.e., RCMs are “driven” by the corresponding GCM. They refine the GCM’s large-scale climate information and yield high-resolution projections [Giorgi2019ThirtyNext]. Due to necessary simplifications and parameterizations, regional climate models differ in their representation of processes, resulting in different RCMs, and initiatives such as the Coordinated Regional Climate Downscaling Experiment (CORDEX) [Giorgi2015RegionalInitiative] provide data from multiple models for several regions around the globe.
@@
However, RCMs remain computationally costly. To address this challenge, statistical downscaling establishes a statistical relationship between low-resolution (LR) and high-resolution (HR) data. In addition to more traditional approaches (e.g. [Maraun2010PrecipitationUser, Maraun2018StatisticalResearch]), machine learning (ML) for resolution enhancement has been applied to predict high-resolution outputs from coarse data in various contexts. This includes downscaling GCMs to RCM resolutions [Rampal2024EnhancingLearning], targeting convection-permitting models [Addison2022] or high-resolution observational data [Harris2022AForecasts, Leinonen2020, PriceStephanRasp2022IncreasingModels, Vosper2023DeepScales], often using varying sets of predictors.
@@
Many different setups for ML-based downscaling have been considered. Among these, some attempts directly approximate the full path from GCM to RCM data: the LR predictors stem from GCM data and the predictands from the RCM [Bano-Medina2021]. We refer to this as RCM emulation. Since there are modeling biases between global and regional outputs, downscaling GCMs to RCMs requires learning a transformation map in addition to the pure resolution refinement [vanderMeer2023DeepFrameworks]. Similar mismatches can arise in other setups, for example with predictors from a biased numerical weather model and targets from radar measurements [Harris2022AForecasts, PriceStephanRasp2022IncreasingModels]. Alternative setups include super-resolution and perfect prognosis downscaling. In the super-resolution setting, the inputs for the ML model are artificially coarsened versions of the HR targets, so that the predictor is exactly a degraded form of the target field [Harder2023Hard-ConstrainedDownscaling, Leinonen2020, Stengel2020, Vosper2023DeepScales]. Other works instead use different climate variables as predictors, e.g. by learning the relationship between large-scale atmospheric predictors and a local HR target. In perfect prognosis downscaling, this relationship is learned from observational data [Bano-Medina2020ConfigurationDownscaling], whereas [Doury2022] and [Doury2024OnPrecipitation] learn it from RCM data. Note that they also refer to this task as “RCM emulation”, which is different from our RCM emulation: they consider both predictors and targets from the RCM, whereas we use predictors from the GCM and targets from the RCM. Some studies conclude that emulators can successfully learn the GCM-RCM map directly [vanderMeer2023DeepFrameworks], whereas others find better performance for the pure super-resolution task compared to a setup that additionally needs to learn large-scale shifts [Harris2022AForecasts]. In addition, learning the transformation map and the downscaling map separately can lead to improved results [PriceStephanRasp2022IncreasingModels].
@@
Many downscaling approaches focus on producing a deterministic map at the fine resolution, given LR predictors [Bano-Medina2020ConfigurationDownscaling, Bano-Medina2021, Doury2022]. However, to capture the full variability, including extreme events, a stochastic approach is needed. In this case, the goal is to model the conditional distribution of the HR target given the LR input. Generative ML methods provide a way to achieve this by “learning” a complex, high-dimensional distribution from training data. After training, generative models can generate samples that reflect the variability present in the training distribution. Examples of generative models are Generative Adversarial Networks (GANs) [Goodfellow2020GenerativeNetworks] and diffusion models [Ho2020DenoisingModels, KarrasElucidatingModels, Song2019GenerativeDistribution, Song2020Score-BasedEquations]. In past studies, both approaches have been used successfull, e.g. to downscale precipitation [Addison2022, Harris2022AForecasts, Leinonen2020, Vosper2023DeepScales] or wind [Miralles2022, Stengel2020, White2019]. While GANs can often generate realistic samples, their optimization can be unstable [Arjovsky2017TowardsNetworks], and they are prone to mode collapse, where the model fails to represent the full diversity of the target distribution. Diffusion models, in contrast, are effective in capturing complex distributions, but are computationally costly, especially during inference.
@@
Besides GANs and diffusion models, it is possible to train a generative model by minimizing a so-called proper scoring rule [Shen2024Engression:Regression, Alet2025SkillfulMarginals, Chen2022GenerativePost-processing, Kochkov2024NeuralClimate, Lang2024AIFS-CRPS:Score, Pacchiardi2024ProbabilisticMinimization]. Such a scoring rule compares the generated distribution by the machine learning model with the target RCM distribution. It is minimal if and only if the generated and target distribution agree. This loss function allows for easy optimization and fast training. In addition, it is flexible and can be combined with different model architectures.
@@
Temporal consistency of downscaled data is important for accurate representation of multi-day statistics, for example in the assessment of heatwaves or accumulated precipitation. Yet, most existing methods do not consider temporal consistency. A few studies perform refinement in a temporal dimension, generating multiple time steps on a high-resolution scale for one or more input fields [Glawion2023SpateGAN:Approach, Serifi2021, Bassetti2024DiffESM:Models]. However, they still process each input patch separately of previous ones and, thus, do not ensure temporal consistency of longer time series on the high-resolution scale. In contrast, AI-based weather forecasting commonly employs an autoregressive approach: previous time steps are included as predictors, and an iterative application of the ML model can provide temporally consistent predictions.
@@
In this study, we emulate the conditional distribution of the RCM data giving the corresponding GCM input. We train our model on a joint dataset combining multiple pairs of RCMs and GCMs. To make best use of this pooled dataset, we introduce a novel approach that splits the problem into several physically meaningful steps, similar to [PriceStephanRasp2022IncreasingModels]. We first correct for large-scale mismatches between the driving GCM and a coarsened version of the target RCM. Second, we perform super-resolution from the coarsened RCM to the high-resolution RCM. For this, we progressively refine the resolution in multiple stages [Shen2025ReverseDistributions]. In addition, to handle the high-dimensional output in a computationally efficient manner, we introduce a new architecture with sparse local stochastic layers. The generated values at a specific location only depend on the values at the lower-resolution in a local neighborhood. This architecture enables learning location-specific features while at the same time limiting the number of model parameters. For each step, we train a generative model minimizing the energy score, a multivariate generalization of the Continuous Ranked Probability Score (CRPS), one example of a proper scoring rule. Accordingly, we call our method EnScale (energy score + downscaling).
@@
    We demonstrate that emulation of RCMs from GCM data in a multivariate and generative setup is feasible, both capturing the variability of RCM targets as well as multivariate dependencies by jointly emulating temperature (tas), precipitation (pr), solar radiation (rsds), and surface wind (sfcWind).
@@
    We propose new methods for RCM emulation: EnScale and EnScale-t, which combine the idea of downscaling via coarse correction with a novel generative model architecture and loss. They achieve great performance, while at the same time being computationally lightweight.
@@
    We establish a comprehensive evaluation framework with metrics for overall performance, spatial and temporal structure, behavior in the extremes, and variable dependencies. For comparison, we propose a set of benchmarks, including physically motivated approaches and state-of-the-art generative models.
@@ 2.1 Downscaling as a generative modeling problem
Figure 1: Illustration of the dataset. The first row shows the GCM data from CNRM-CM5 on two example days with similar precipitation fields, displayed on the full spatial extent that is used as the model input. The second row presents the same GCM data cropped to approximately match the target area used in this study. The fourth row shows the corresponding target RCM data (ALADIN63 and RegCM4-6) that downscaled this GCM. The third row presents these RCM fields manually coarsened to a resolution comparable to the GCM data. Small differences between shape and domain of cropped GCM (row two) and coarsened RCM (row three) are due to small resolution differences and misaligned grids.
@@ 2.2 Downscaling with coarse correction
Our model is trained jointly on the entire dataset comprising multiple pairs of GCM data $X$ and corresponding RCM data $Y$. They stem from three different GCMs, and each driving GCM was downscaled by several RCMs (between two and three RCMs for one GCM, see Sec. 3). Despite training a single unified model, our goal is to capture both GCM- and RCM-specific characteristics. We aim to learn the conditional $p^{\iota}_{Y|X}$. Here, $\iota$ is an index for the GCM-RCM pair, emphasizing that we target the distribution for each GCM-RCM pair individually. Fig. 1 shows example GCM data on two selected days as well as corresponding downscaled data from two RCMs. These examples help to illustrate several challenges present in our heterogeneous dataset, as explained in the following.
@@
Secondly, each GCM-RCM pair follows a different distribution, and the ML model needs to learn the correct downscaling function for each of them. Fig. 1 compares the downscaled fields by two different RCMs, given the same driving GCM. We find that responses from multiple RCMs to the same GCM input differ considerably. Partly, this can be due to internal variability, as discussed above. In addition, there are disagreements between RCMs. This is because each RCM has parameterizations describing the effect of small-scale processes (e.g., clouds) that differ from those of other RCMs and the driving GCM. The land surface may be fixed in one model and respond to the atmosphere in another. Our experiments suggest that most differences between the RCMs are already present in the first step from GCM to coarsened RCM data, see App. A.1.
@@
To solve these challenges, we split the problem into two parts, separating the correction on coarse scales and the super-resolution task [PriceStephanRasp2022IncreasingModels, Wan2024StatisticalModels], as visualized in Fig. 2. In the first step, we train a generative model to learn the conditional distribution of coarsened RCM data $Z$ given the GCM inputs $X$, $p^{\iota}_{Z|X}$. For this, the RCM data is artificially coarsened through average pooling over patches of size $16\times 16$, producing additional data $Z$ for each day. As the resolutions of $X$ ($\approx$250\text{\,}\mathrm{k}\mathrm{m}$\$) and $Z$ ($\approx$170\text{\,}\mathrm{k}\mathrm{m}$\$) are similar, predicting $Z$ from $X$ does not include any “downscaling”. Instead, this step captures large-scale mismatches between GCM and RCM. The output of this step is relatively low-dimensional (dimension $8\times 8$ per variable) and we call this the coarse model.
@@
For the second step, we learn the conditional distribution $p^{\iota}_{Y|Z}$. This is a pure super-resolution problem, as $Z$ is the average-pooled version of $Y$, excluding the large-scale mismatches between $X$ and $Y$. Hence, we refer to the corresponding generative model as the super-resolution model. The transformation from $Z$ to $Y$ is high-dimensional, but the training data is more homogeneous as the different GCM-RCM pairs behave very similarly in this step (see App. A.1). We split the super-resolution model again into several resolution steps, each one learning a simpler conditional distribution (see Sec. 2.4).
In each step of our modeling chain, we add predictor variables indicating the GCM and RCM considered (see App. B.4 for how these predictors are integrated in the model architectures). This allows our model to learn the conditional $p^{\iota}_{Y|X}$ for each pair individually. It enables capturing model-specific details, while at the same time leveraging shared properties from all pairs, benefiting from a large dataset.
@@
All steps are trained independently. We tested more complex training procedures but found no improvement in performance. In addition, each step uses the variables defined for that stage, for example, the super-resolution model only uses $Z$ as input and not also $X$. This simplification approximates the full distribution $p^{\iota}_{Y|X}$ accurately if $Z$ contains all the relevant information about the target $Y$ or, in other words, the map from $Z$ to $Y$ does not depend on $X$ anymore. This approximation is formalized through conditional independence and checked heuristically in App. B.1. Empirically, we found no (e.g., precipitation) or at most 2 % (e.g., temperature) decrease in performance due to this approximation.

### v3

@@ Abstract
The practical use of future climate projections from global circulation models (GCMs) is often limited by their coarse spatial resolution, requiring downscaling to generate high-resolution data. Regional climate models (RCMs) provide this refinement, but are computationally expensive. To address this issue, machine learning (ML) models can learn the downscaling function, mapping coarse GCM outputs to high-resolution fields. Among these, generative approaches aim to capture the full conditional distribution of RCM data given coarse-scale GCM data, which is characterized by large variability and thus challenging to model accurately. We introduce EnScale, a generative ML framework emulating the full GCM-to-RCM map by training on multiple pairs of GCM and corresponding RCM data. It first adjusts large-scale mismatches between GCM and coarsened RCM data, followed by a super-resolution step to generate high-resolution fields. To efficiently model the high-dimensional output, the super-resolution step employs a novel class of sparse local stochastic layers. Both steps employ generative models optimized with the energy score, a proper scoring rule. Compared to state-of-the-art ML downscaling approaches, our setup reduces computational cost by about one order of magnitude. EnScale jointly emulates multiple variables – temperature, precipitation, solar radiation, and wind – spatially consistent over Central Europe. In addition, we propose a variant EnScale-t that enables temporally consistent downscaling. We establish a comprehensive evaluation framework across various categories including calibration, spatial and temporal structure, extremes, and multivariate dependencies. Comparison with diverse benchmarks demonstrates EnScale(-t)’s competitive performance and computational efficiency, offering a promising approach for accurate and temporally consistent RCM emulation.
@@ 1 Introduction
General circulation models (GCMs) provide us with future climate projections that enable assessing potential changes in climate under different greenhouse gas emission scenarios. Multiple research groups around the world develop and maintain GCMs that incorporate different modeling assumptions and parameterizations to simulate the climate system of the Earth. By virtue of international efforts such as the Coupled Model Intercomparison Project (CMIP) [16, 68], large ensembles of GCM data are available for various Representative Concentration Pathway (RCP) / Shared Socioeconomic Pathway (SSP) scenarios.
@@
For computational reasons, these physics-based models are typically operated at a coarse spatial resolution. GCMs are able to simulate large-scale climate patterns, but often lack the spatial resolution needed to capture climate variations at impact-relevant regional and local scales [8, 43, 51, 78]. Downscaling describes the process of increasing the spatial and/or temporal resolution of GCM outputs. The common approach of dynamical downscaling relies on process-based regional climate models (RCMs) that use GCM data as boundary conditions, i.e., RCMs are “driven” by the corresponding GCM. They refine the GCM’s large-scale climate information and yield high-resolution projections [18]. Due to necessary simplifications and parameterizations, regional climate models differ in their representation of processes, resulting in different RCMs, and initiatives such as the Coordinated Regional Climate Downscaling Experiment (CORDEX) [17] provide data from multiple models for several regions around the globe.
@@
However, RCMs remain computationally costly. To address this challenge, statistical downscaling establishes a statistical relationship between low-resolution (LR) and high-resolution (HR) data. In addition to more traditional approaches (e.g. [40, 42]), machine learning (ML) for resolution enhancement has been applied to predict high-resolution outputs from coarse data in various contexts. This includes downscaling GCMs to RCM resolutions [51], targeting convection-permitting models [1] or high-resolution observational data [25, 37, 48, 71], often using varying sets of predictors.
@@
Many different setups for ML-based downscaling have been considered. Among these, some attempts directly approximate the full path from GCM to RCM data: the LR predictors stem from GCM data and the predictands from the RCM [6]. We refer to this as RCM emulation. Since there are modeling biases between global and regional outputs, downscaling GCMs to RCMs requires learning a transformation map in addition to the pure resolution refinement [69]. Similar mismatches can arise in other setups, for example with predictors from a biased numerical weather model and targets from radar measurements [25, 48]. Alternative setups include super-resolution and perfect prognosis downscaling. In the super-resolution setting, the inputs for the ML model are artificially coarsened versions of the HR targets, so that the predictor is exactly a degraded form of the target field [24, 37, 65, 71]. Other works instead use different climate variables as predictors, e.g. by learning the relationship between large-scale atmospheric predictors and a local HR target. In perfect prognosis downscaling, this relationship is learned from observational data [5], whereas [13] and [14] learn it from RCM data. Note that they also refer to this task as “RCM emulation”, which is different from our RCM emulation: they consider both predictors and targets from the RCM, whereas we use predictors from the GCM and targets from the RCM. The literature shows mixed results: Some studies conclude that emulators can successfully learn the GCM-RCM map directly [69], whereas others find better performance for the pure super-resolution task compared to a setup that additionally needs to learn large-scale shifts [25], and some highlight the potential of learning the transformation map and the downscaling map separately [48].
@@
Many downscaling approaches focus on producing a deterministic map at the fine resolution, given LR predictors [5, 6, 13]. However, to capture the full variability, including extreme events, a stochastic approach is needed. In this case, the goal is to model the conditional distribution of the HR target given the LR input. Generative ML methods provide a way to achieve this by “learning” a complex, high-dimensional distribution from training data. After training, generative models can generate samples that reflect the variability present in the training distribution. Examples of generative models are Generative Adversarial Networks (GANs) [23] and diffusion models [29, 33, 63, 64]. In past studies, both approaches have been used successfully, e.g. to downscale precipitation [1, 25, 37, 71] or wind [45, 65, 77]. While GANs can often generate realistic samples, their optimization can be unstable [3], and they are prone to mode collapse, where the model fails to represent the full diversity of the target distribution. Diffusion models, in contrast, are effective in capturing complex distributions, but are computationally costly, especially during inference. Reducing computational latency is an active area of research, e.g., via efficient sampling algorithms [61, 39], latent diffusion models [53], and distillation methods [55, 62]. While these approaches can accelerate inference, they may introduce trade-offs in sample quality depending on compression levels or the number of denoising steps.
@@
Besides GANs and diffusion models, it is possible to train a generative model by minimizing a proper scoring rule [60, 2, 11, 34, 36, 46]. Such a scoring rule compares the generated distribution by the machine learning model with the target RCM distribution. It is minimal if and only if the generated and target distribution agree [22]. This loss function allows for easy optimization and fast training. In addition, it is flexible and can be combined with different model architectures. While the Continuous Ranked Probability Score (CRPS) recently gained attention in AI-based weather forecasting, its multivariate generalization, the energy score, has been adopted as a training objective only in a few recent studies [59, 58, 46] and remains relatively underexplored. To our knowledge, training via proper scoring rules has not yet been rigorously investigated the case of climate downscaling.
@@
Temporal consistency of downscaled data is important for accurate representation of multi-day statistics, for example in the assessment of heatwaves or accumulated precipitation. Yet, most existing methods do not consider temporal consistency. A few studies perform refinement in a temporal dimension, generating multiple time steps on a high-resolution scale for one or more input fields [19, 57, 7]. However, they still process each input patch separately of previous ones and, thus, do not ensure temporal consistency of longer time series on the high-resolution scale. In contrast, AI-based weather forecasting commonly employs an autoregressive approach: previous time steps are included as predictors, and an iterative application of the ML model can provide temporally consistent predictions.
@@
In this study, we emulate the conditional distribution of the RCM data giving the corresponding GCM input for a domain covering Central Europe. We train our model on a joint dataset combining multiple pairs of RCMs and GCMs. To make best use of this pooled dataset, we introduce a novel approach that splits the problem into several physically meaningful steps, similar to [48] and [38]. We first correct for large-scale mismatches between the driving GCM and a coarsened version of the target RCM. Second, we perform super-resolution from the coarsened RCM to the high-resolution RCM. For this, we progressively refine the resolution in multiple stages [58]. In addition, to handle the high-dimensional output in a computationally efficient manner, we introduce a new architecture with sparse local stochastic layers. The generated values at a specific location only depend on the values at the lower-resolution in a local neighborhood. This architecture enables learning location-specific features while at the same time limiting the number of model parameters. For each step, we train a generative model minimizing the energy score. Accordingly, we call our method EnScale (energy score + downscaling).
@@
    We demonstrate that fully data-driven emulation of RCMs from GCM data in a multivariate and generative setup is feasible, both capturing the variability of RCM targets as well as multivariate dependencies by jointly emulating temperature (tas), precipitation (pr), solar radiation (rsds), and surface wind (sfcWind).
@@
    We propose new methods for RCM emulation: EnScale and EnScale-t, which combine the idea of downscaling via coarse correction with a novel generative model architecture, optimized for using the energy score as a loss function. This loss allows accurate representation of the target distribution, while at the same time yielding computationally lightweight models.
@@
    We establish a comprehensive evaluation framework with metrics for overall performance, spatial and temporal structure, behavior in the extremes, variable dependencies and climate change signal. For comparison, we propose a set of benchmarks, including physically motivated approaches and state-of-the-art generative models.
@@ 2.1 Downscaling as a generative modeling problem
Figure 1: Illustration of the dataset. This highlights the difference between RCMs, the mismatches of GCM and RCM even at a coarse scale, and the large variability of RCM fields given GCM conditions. The first row shows the GCM data from CNRM-CM5 on two example days with similar precipitation fields, displayed on the full spatial extent that is used as the model input. The second row presents the same GCM data cropped to approximately match the target area used in this study. The fourth row shows the corresponding target RCM data (ALADIN63 and RegCM4-6) that downscaled this GCM. The third row presents these RCM fields manually coarsened to a resolution comparable to the GCM data. Small differences between shape and domain of cropped GCM (row two) and coarsened RCM (row three) are due to small resolution differences and misaligned grids.
@@ 2.2 Downscaling with coarse correction
Our model is trained jointly on the entire dataset comprising multiple pairs of GCM data $X$ and corresponding RCM data $Y$. They stem from three different GCMs, and each driving GCM was downscaled by several RCMs (between two and three RCMs for one GCM, see Sec. 3). Despite training a single unified model, our goal is to capture RCM-specific characteristics. We aim to learn the conditional $p^{\iota}_{Y|X}$. Here, $\iota$ is an index for the RCM, emphasizing that we target the distribution for each RCM individually. Fig. 1 shows example GCM data on two selected days as well as corresponding downscaled data from two RCMs. These examples help to illustrate several challenges present in our heterogeneous dataset, as explained in the following.
@@
Secondly, each GCM-RCM pair follows a different distribution. While we train the ML model jointly with all pairs to maximize the available dataset size, we still aim for the ML model to learn the correct downscaling function for each pair individually by conditioning on the model on the RCMs considered. We also tested conditioning on the GCM information. However, we found that this does not improve the result. Hence, to allow our model to be applicable to unseen GCMs, we decided to only add a label for the RCM. Fig. 1 compares the downscaled fields by two different RCMs, given the same driving GCM. We find that responses from multiple RCMs to the same GCM input differ considerably. Partly, this can be due to internal variability, as discussed above. In addition, there are disagreements between RCMs. This is because each RCM has parameterizations describing the effect of small-scale processes (e.g., clouds) that differ from those of other RCMs and the driving GCM [18]. The land surface may be fixed in one model and respond to the atmosphere in another.
@@
To solve these challenges, we split the problem into two parts, separating the correction on coarse scales and the super-resolution task [48, 74], as visualized in Fig. 2. In the first step, we train a generative model to learn the conditional distribution of coarsened RCM data $Z$ given the GCM inputs $X$, $p^{\iota}_{Z|X}$. For this, the RCM data is artificially coarsened through average pooling over patches of size $16\times 16$, producing additional data $Z$ for each day. As the resolutions of $X$ ($\approx$250\text{\,}\mathrm{k}\mathrm{m}$\$) and $Z$ ($\approx$170\text{\,}\mathrm{k}\mathrm{m}$\$) are similar, predicting $Z$ from $X$ does not include any “downscaling”. Instead, this step captures large-scale mismatches between GCM and RCM. The output of this step is relatively low-dimensional (dimension $8\times 8$ per variable) and we call this the coarse model.
@@
For the second step, we learn the conditional distribution of high-resolution RCM data $Y$ given coarsened RCM data $Z$, $p^{\iota}_{Y|Z}$. This is a pure super-resolution problem, as $Z$ is the average-pooled version of $Y$, excluding the large-scale mismatches between $X$ and $Y$. Hence, we refer to the corresponding generative model as the super-resolution model. The transformation from $Z$ to $Y$ is high-dimensional, but the training data is more homogeneous across GCM-RCM pairs: To assess this, we tested training a single joint model for all pairs without conditioning on the model identifier. We found that this primarily affects the loss for the coarse model, while we only observe marginal performance decrease for the super-resolution model. This suggests that different GCM-RCM pairs behave similarly in the super-resolution step.
We split the super-resolution model again into several resolution steps, each one learning a simpler conditional distribution [58]. In particular, each step increases resolution by a factor of two in both spatial dimensions, i.e., we model $p^{\iota}_{Z^{(k+1)}|Z^{(k)}}$, where $Z^{(k)},Z^{(k+1)}$ are average-pooled versions of the target $Y$ with dimensions $2^{k}\times 2^{k}$ and $2^{k+1}\times 2^{k+1}$, respectively. The first step uses $k=3$, i.e. $Z=Z^{(3)}$, corresponding to the dimension of $Z$ ($8\times 8$). Then, $k$ increases by 1 for each step, so that the last step uses $k=6$, as $Y$ has dimension $128\times 128$.
In each step of our modeling chain, we train jointly on the entire dataset, but add predictor variables indicating the RCM considered (see App. C.4 for how these predictors are integrated in the model architectures).
@@
All steps are trained independently. We tested more complex training procedures but found no improvement in performance. In addition, each step uses the variables defined for that stage, for example, the super-resolution model only uses the coarsened RCM data $Z$ as input and not also the GCM data $X$. This simplification approximates the full distribution $p^{\iota}_{Y|X}$ accurately if $Z$ contains all the relevant information about the target $Y$ or, in other words, the map from $Z$ to $Y$ does not depend on $X$ anymore. This approximation is formalized through conditional independence and checked heuristically in App. C.1. Empirically, we found no (e.g., precipitation) or at most 2 % (e.g., temperature) decrease in performance due to this approximation.